The question says
What amount of current will pass through 100 watt coil if it takes 21 seconds for the coil to just melt a lump of ice of mass 10g at 0
1
Expert's answer
2021-07-07T08:38:52-0400

We can find the energy needed to melt the lump of ice as follows:


"E=Pt,"

here, "P=I^2R" is the power dissipated by the coil, "I" is the amount of current passed through the coil, "R=10\ \Omega" is the resistance of the coil, "t=21\ s" is the time needed to melt the lump of ice.

From the other hand, we can write:


"E=mL_f,"

here, "m=0.01\ kg" is the mass of the ice, "L_f=3.36\cdot10^5\ \dfrac{J}{kg}."

Finally, we can equate two expressions and find the amount of current passed through the coil:


"I^2Rt=mL_f,""I=\sqrt{\dfrac{mL_f}{Rt}}=\sqrt{\dfrac{0.01\ kg\cdot3.36\cdot10^5\ \dfrac{J}{kg}}{10\ \Omega\cdot21\ s}}=4\ A."